OpenStudy (anonymous):

I'd like to differentiate log 10^(sq root t). Thanks

zepdrix (zepdrix):

\[\Large\rm y=\log \sqrt t\]We need to find a derivative? Do you recall your derivative for natural log?

zepdrix (zepdrix):

We can apply a change of base,\[\Large\rm \log_{a}(b)=\frac{\ln(b)}{\ln(a)}\]and then apply what we know with the natural log derivative.

zepdrix (zepdrix):

\[\Large\rm y=\log_{10} \sqrt t\] \[\Large\rm y=\frac{\ln \sqrt t}{\ln(10)}\]No need for quotient rule or anything fancy like that. The denominator is just constant, let's pull it out of the way.\[\Large\rm y=\left(\frac{1}{\ln(10)}\right)\ln \sqrt t\]

zepdrix (zepdrix):

Where you at Larry? What are you thinking? :)

OpenStudy (anonymous):

No need to change base at all. Simply the quest can be written as: 0.5logt. Then derivative is: 0.5/(t.ln10) That's it.

zepdrix (zepdrix):

Change of base is better, that's where the 1/(ln10) is coming from anyway. You skipped a step by memorizing some weird shortcut.